OpenStudy (anonymous):

what is the slope between the two points (2, 3) and (1, 4)

OpenStudy (anonymous):

It's the change in the y value divided by the change in the x value.

OpenStudy (anonymous):

slope = y1 - y2 / x1-x2 so it is slope = 4-3 / 1-2 slope = 1/-1 = -1

OpenStudy (karatechopper):

y2-y1/x2-x1

OpenStudy (anonymous):

So the slope would be -1?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Can I ask a couple more of you?

OpenStudy (anonymous):

yea post em up

OpenStudy (anonymous):

Notice that although @jayz657 and @karatechopper have different formulas (the order is different), they both product the same result. In other words, it doesn't matter whether you subtract the second point values from the first or vice-versa, as long as you don't mix the order (y2 - y1)/(x1 -x2) would be wrong.

OpenStudy (anonymous):

Okay. First is what is the slope between the two points (–3, 1) and (3, –6) and the other one isonsider the line that passes through the point (–3, 2) and has a slope of 3. Part 1: Write the equation of this line using point-slope form. (2 points) Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. (2 points) Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work. (2 points) (6 points)

OpenStudy (anonymous):

It's on a practice test so I need some help with them.

OpenStudy (anonymous):

Thanks in advance

OpenStudy (anonymous):

so using the same equation as before slope = -6 -1 / 3-(-3) = -7/6 as your slope for the next part remember the equation of a line is in slope intercept form is y=mx+b , where m=slope, b = y- intercept so your given m =3 and (x,y) = (-3,2) so plug everything in and solve for b 2 = 3(-3) + b 2 = -9 + b b = 7 so your answer is y=3x+7

OpenStudy (anonymous):

Uh would that be 6/7 -6/7 7/ or -7/6?

OpenStudy (anonymous):

and for standard you need to put it in the form: mx + y = b so it would be -3x + y = 7 <- might need someone to double check this i havent done this in a long time

OpenStudy (anonymous):

its -7/6

OpenStudy (anonymous):

Okay. and the Second one?

OpenStudy (anonymous):

look at my other post i made its all there

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

np glad to help